# Rephrase and Contrast: Fine-Tuning Language Models for Enhanced Understanding of Communication and Computer Networks

### v1

@@ Abstract
For data mining, we leverage the language processing ability of GPT-4 to analyze raw materials extracted from classic networking textbooks and generate high-quality QA pairs that meet the needs of RaC fine tuning. The QA pair generation process applies the in-context learning (ICL) technique to guide the GPT-4 model in structuring the data and help the model better understand what output is desired in each step.
@@ II Related Work
Methodologies: Most previous works were built upon prompt tuning over off-the-shelf models with only a few considered model fine-tuning. For example, in [8], the authors put forth a framework that empowers LLMs with various model-optimizing techniques including model fine-tuning. However, the work only conducted experiments on prompt-tuning methods. In [2, 4, 5, 9, 10, 11, 8, 6], few-shot prompting methods, such as in-context learning (ICL), are used to guide LLMs to generate answers reliably. Chain of thought (CoT) were applied in [2, 10, 6, 8] to enhance LLMs’ reasoning ability. Additionally, retrieval methods like retrieval augmented generation (RAG) are used in [13, 8] to enhance the off-the-shelf LLM’s knowledge in wireless networks.
@@ IV Training and Testing Datasets
The extensive dataset required for LLM fine-tuning necessitates an automated approach for QA pair generation. We leverage the GPT-4 model to generate a substantial corpus of high-quality QA pairs. Importantly, we highlight that the QA pair we are looking for is more than a simple combination of a question and an answer. A core feature of our QA pair is the the inclusion of problem rephrasing and contrasting statements, which are required in the RaC framework. The rest of this paper refers to the new QA pair as the RaC QA pair to distinguish with the conventional QA pair.
@@ IV-A Dataset Construction
ChoiceBoost is designed to mitigate the above two issues. In ChoiceBoost, each question in the dataset was duplicated four times, with each duplicate assigned a different correct answer (A, B, C, or D), while maintaining the original order of incorrect options. This approach substantially expands the dataset without additional API cost (which enhances the model’s capacity by learning from diverse answer configurations) and mitigating biases inherent in the initial data generation process.
@@ IV-B Data Resources Released
The initial dataset derived from ten textbooks without post-processing data augmentation comprises 15,119 QA pairs. These original QA pairs encompass essential sub-domains knowledge of computer networking described in each individual textbook, such as layered network architecture, protocols design, network security, and network management. From the raw data, 756 QA pairs (approximately 5%) were randomly selected to form a testing problem set. We then used ChoiceBoost to argument the remaining 15,119 – 756 = 14,363 QA pairs, resulting in an unbiased training dataset containing 14,363 * 4 = 57,452 QA pairs.
@@
In addition to the GPT-generated QA pairs, we manually collected 350 multiple-choice testing questions closely aligned with the content of the ten selected textbooks. These questions were sourced from publicly available resources, including open-source exam papers in online databases and problem sets provided within the textbooks. Note that these QA pairs are exclusively used for testing purposes only, and they lack problem rephrasing and answer explanations required in the RaC framework. The value of the new testing problem set lies in their focus on complex logical reasoning and intricate calculations. For instance, such problem might require calculating the maximum network throughput under changing bandwidth conditions, or involve multi-step reasoning, such as optimizing a routing algorithm considering traffic fluctuations and delay constraints. The motivation for building the more challenging dataset stems from our observation that GPT-generated QA pairs may not adequately address complex reasoning and mathematical computations due to the limitations of the GPT-4 model in these areas. Instead, the GPT-generated pairs primarily emphasize concept comprehension. For example, a typical problem from the set might ask, ’What is the function of a router in a network?’ or ’Which OSI layer is responsible for error detection and correction?’. While the 756 GPT-generated testing QA pairs can assess a model’s grasp of networking concepts, a testing problem set concentrating on logical reasoning and mathematical computation is equally essential for a comprehensive evaluation of the fine-tuned model. For the rest of this paper, we refer to the testing problem set containing 756 GPT-generated problems as the ”easy problem set” and the 350 manually collected problems as the ”hard problem set.”
@@
Fig. 3 visually illustrates the data distribution of the training dataset and the easy/hard problem set. In this figure, networking knowledge is categorized into nine distinct sub-domains. The figure shows that, the training data, the easy test problem sets, and the hard test problems, all exhibit a balanced coverage of these sub-domains, thereby ensuring that the model is trained and evaluated with a comprehensive and generalized focus on the networking field. These two word clouds further detail the sub-domain content distribution. The left word cloud highlights the foundational and all-inclusive topics in the training dataset and easy problem set; the right one, on the other hand, reflects hard problem set focuses on more complex and detailed aspects of networking.
@@
The easy and hard problem sets constitute two critical testing benchmarks that assess an LLM’s performance in the networking domain from contrasting perspectives. An additional crucial aspect of the model’s evaluation is its performance in a comprehensive and practical problem set with both easy and hard problems. To this end, we randomly down-sampled the easy problem set to create a subset comparable in size to the hard dataset (i.e., 350 QA pairs). Subsequently, we merged this new subset with the hard problem set to form a comprehensive dataset. All three problem sets (easy, hard, and comprehensive) are used in Section V for model evaluation.
@@ V Experiments
For evaluation results on the easy testing problem set, our model correctly answers 80% of questions, significantly surpassing all other baseline models in terms of accuracy. The original LLaMA-2 7B model (without fine-tuning) performs the poorest among all tested models, with only 48% accuracy on the tested dataset. The huge accuracy gap between our fine-tuned model and the foundational LLaMA-2 7B indicates the effectiveness of our method when applied on the LLaMA-2 7B model. Other baseline models of approximately 7B parameter size, such as ChatGLM and BaiChuan, also demonstrate better performance than LLaMA-2 7B, but still fall short of our model, with accuracies ranging from 66% to 70%.

### v2

@@ Abstract
For data mining, we leverage the language processing ability of GPT-4 to analyze raw materials extracted from classic networking textbooks and generate high-quality QA pairs that meet the needs of RaC fine-tuning. The QA pair generation process applies the in-context learning (ICL) technique to guide the GPT-4 model in structuring the data and help the model better understand what output is desired in each step.
@@ II Related Work
Methodologies: Most previous works were built upon prompt tuning over off-the-shelf models with only a few considered model fine-tuning. For example, in [8], the authors put forth a framework that empowers LLMs with various model-optimizing techniques including model fine-tuning. However, the work only conducted experiments on prompt-tuning methods. In [2, 4, 5, 9, 10, 11, 8, 6], few-shot prompting methods, such as in-context learning (ICL), are used to guide LLMs to generate answers reliably. Chain of thought (CoT) was applied in [2, 10, 6, 8] to enhance LLMs’ reasoning ability. Additionally, retrieval methods like retrieval augmented generation (RAG) are used in [13, 8] to enhance the off-the-shelf LLM’s knowledge in wireless networks.
@@ IV Training and Testing Datasets
The extensive dataset required for LLM fine-tuning necessitates an automated approach for QA pair generation. We leverage the GPT-4 model to generate a substantial corpus of high-quality QA pairs. Importantly, we highlight that the QA pair we are looking for is more than a simple combination of a question and an answer. A core feature of our QA pair is the inclusion of problem rephrasing and contrasting statements, which are required in the RaC framework. The rest of this paper refers to the new QA pair as the RaC QA pair to distinguish it from the conventional QA pair.
@@ IV-A Dataset Construction
ChoiceBoost is designed to mitigate the above two issues. In ChoiceBoost, each question in the dataset was duplicated four times, with each duplicate assigned a different correct answer (A, B, C, or D), while maintaining the original order of incorrect options. This approach substantially expands the dataset without additional API cost (which enhances the model’s capacity by learning from diverse answer configurations) and mitigates biases inherent in the initial data generation process.
@@ IV-B Data Resources Released
The initial dataset derived from ten textbooks without post-processing data augmentation comprises 15,119 QA pairs. These original QA pairs encompass essential sub-domains knowledge of computer networking described in each individual textbook, such as layered network architecture, protocols design, network security, and network management. From the raw data, 756 QA pairs (approximately 5%) were randomly selected to form a testing problem set. We then used ChoiceBoost to augment the remaining 15,119 – 756 = 14,363 QA pairs, resulting in an unbiased training dataset containing 14,363 * 4 = 57,452 QA pairs.
@@
In addition to the GPT-generated QA pairs, we manually collected 327 multiple-choice testing questions closely aligned with the content of the ten selected textbooks. These questions were sourced from publicly available resources, including open-source exam papers in online databases and problem sets provided within the textbooks. Note that these QA pairs are exclusively used for testing purposes only, and they lack problem rephrasing and answer explanations required in the RaC framework. The value of the new testing problem set lies in its focus on complex logical reasoning and intricate calculations. For instance, such problems might require calculating the maximum network throughput under changing bandwidth conditions, or involve multi-step reasoning, such as optimizing a routing algorithm considering traffic fluctuations and delay constraints. The motivation for building the more challenging dataset stems from our observation that GPT-generated QA pairs may not adequately address complex reasoning and mathematical computations due to the limitations of the GPT-4 model in these areas. Instead, the GPT-generated pairs primarily emphasize concept comprehension. For example, a typical problem from the set might ask, ’What is the function of a router in a network?’ or ’Which OSI layer is responsible for error detection and correction?’. While the 756 GPT-generated testing QA pairs can assess a model’s grasp of networking concepts, a testing problem set concentrating on logical reasoning and mathematical computation is equally essential for a comprehensive evaluation of the fine-tuned model. For the rest of this paper, we refer to the testing problem set containing 756 GPT-generated problems as the ”easy problem set” and the 327 manually collected problems as the ”hard problem set.”
@@
Fig. 3 visually illustrates the data distribution of the training dataset and the easy/hard problem set. In this figure, networking knowledge is categorized into nine distinct sub-domains. The figure shows that, the training data, the easy test problem sets, and the hard test problems, all exhibit a balanced coverage of these sub-domains, thereby ensuring that the model is trained and evaluated with a comprehensive and generalized focus on the networking field. These two word clouds further detail the sub-domain content distribution. The left word cloud highlights the foundational and all-inclusive topics in the training dataset and easy problem set; the right one, on the other hand, reflects the hard problem set focuses on more complex and detailed aspects of networking.
@@
The easy and hard problem sets constitute two critical testing benchmarks that assess an LLM’s performance in the networking domain from contrasting perspectives. An additional crucial aspect of the model’s evaluation is its performance in a comprehensive and practical problem set with both easy and hard problems. To this end, we randomly down-sampled the easy problem set to create a subset comparable in size to the hard dataset (i.e., 327 QA pairs). Subsequently, we merged this new subset with the hard problem set to form a comprehensive dataset. All three problem sets (easy, hard, and comprehensive) are used in Section V for model evaluation.
@@ V Experiments
For evaluation results on the easy testing problem set, our model correctly answers 80% of questions, significantly surpassing all other baseline models in terms of accuracy. The original LLaMA-2 7B model (without fine-tuning) performs the poorest among all tested models, with only 48% accuracy on the tested dataset. The huge accuracy gap between our fine-tuned model and the foundational LLaMA-2 7B indicates the effectiveness of our method when applied to the LLaMA-2 7B model. Other baseline models of approximately 7B parameter size, such as ChatGLM and BaiChuan, also demonstrate better performance than LLaMA-2 7B, but still fall short of our model, with accuracies ranging from 66% to 70%.